OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form

OpenStudy (luigi0210):

distribute and then move terms all to the right leaving it as y=mx+b

OpenStudy (luigi0210):

*Ax+By=C

OpenStudy (luigi0210):

That's slope-intercept form.. I think standard form is AX+By=C

OpenStudy (jhannybean):

Oh is it... haha, i see. ok.

OpenStudy (luigi0210):

Aw, why'd you delete your work?

OpenStudy (jhannybean):

Because it's the wrong explanation lol. I just woke up -_-

OpenStudy (luigi0210):

Haha, doesn't mean it wasn't helpful. So yea @MakeMeSmile332 just do that and let's see what you get :)

OpenStudy (anonymous):

4x - y = -22?

OpenStudy (luigi0210):

What are you typing Jhann, a novel? :P and correct

OpenStudy (jhannybean):

Standard form \[\large \color{green}Ax+\color{blue}By=C\]you have your y's and x's on both sides of the equation, so you want to put them all on one side, and put all your numbers on the other. \[\large \color{blue}y-2=4(\color{green}x+5)\]start by distributing the 4 to (x+5)\[\large \color{blue}y-2=4(\color{green}x)+4(5)\]now subtract your 4(x) so it can be on the left side. \[\large \color{blue}y-2-4(\color{green}x)=-4(\color{green}{x})+4(\color{green}{x})+4(5)\] add +2 to both sides to get rid of the 2 on the left side. \[\large \color{blue}y-2+2-4(\color{green}x)=4(5)+2\]simplify \[\large \color{blue}y-4(\color{green}x)=4(5)+2\]

OpenStudy (jhannybean):

haha. good job

OpenStudy (jhannybean):

It'll be +22, :P not -22!

OpenStudy (luigi0210):

We got -22

OpenStudy (jhannybean):

really? O_o

OpenStudy (luigi0210):

\[y-2=4(x+5)\] \[y-2=4x+20\] \[-4x+y=22\] multiply a negative since Ax cannot be negative \[4x-y=-22\]

OpenStudy (jhannybean):

Oh, I just looked it up! you CAN have a negative Ax, but it's just proper to multiply a negative across to fit the standard form of the equation better

OpenStudy (luigi0210):

Ha, well then. @MakeMeSmile332 nice job, you didn't need us at all :P

OpenStudy (jhannybean):

xD goood job!